5/8 - 2/5 = 1 over Y. How do i solve for Y
Answer by Alan3354(69443)

5/8 - 2/5 = 1 over Y
5/8 - 2/5 = 1/y
25/40 - 16/40 = 1/y
9/40 = 1/y
y = 40/9